OpenStudy (anonymous):

What is the solution to the equation? 7(5 + 9) = 7 · x + 7 · 9 A. x = 1 B. x = 5 C. x = 7 D. x = 9

OpenStudy (hitaro9):

So first you're always going to want to do stuff that's inside the parenthesis first So 7(5+9) = 7(14)

OpenStudy (raffle_snaffle):

add then multi.

OpenStudy (hitaro9):

Now, there's nothing else on that side of the equation, so multiplying it out is pretty easy. 98 = 7 · x + 7 · 9

OpenStudy (raffle_snaffle):

@Hitaro9 she doesn't want to solve the problem... She is expecting us to just give her the dang answer...

OpenStudy (anonymous):

No, I'm listening. Don't judge a book by its cover sir. >_<

OpenStudy (raffle_snaffle):

I don't have to judge...=P

OpenStudy (hitaro9):

@Akayla do you actually have to do the dot product or just multiply?

OpenStudy (anonymous):

the dot is just another way of saying "multiplication" without doing the typical multiplication x